OpenStudy (anonymous):

Another one is x/(2x-1)^1/2 with an upper interval of 5 and the lower being 1

myininaya (myininaya):

try letting u=2x-1 so du=2dx

myininaya (myininaya):

since you have x on top, you will also need to put that in terms of u. since u=2x-1, then x=(u+1)/2. The upper limit is 2(5)-1 and the lower limit is 2(1)-1

myininaya (myininaya):

\[1/4\int\limits_{1}^{9}(u+1)/(u ^{1/2}) du\]

myininaya (myininaya):

\[1/4 \int\limits_{1}^{9}(u ^{1/2}+u^{-1/2} )du\]

OpenStudy (anonymous):

Where are getting 1/4 from?